write an equation for a line through the points (3,6) and (-6,-9)
Answer by Nate(3500)

Linear Line:
(3,6) and (-6,-9)
m = (y2 - y1)/(x2 - x1)
m = (6 + 9)/(3 + 6) = 15/9 = 5/3
y = m(x - x1) + y1
y = (5/3)(x - 3) + 6
y = (5/3)x - 5 + 6
y = (5/3)x + 1